OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (4, 3) and a directix of y = 13? f(x) = −one twentieth (x − 4)2 + 10 f(x) = −one twentieth (x − 4)2 + 8 f(x) = one twentieth (x − 4)2 + 10 f(x) = one twentieth (x + 4)2 + 8

OpenStudy (campbell_st):

the focus is below the directrix so its concave down...  the distance between the focus and directrix is 2* focal length. so disstance is 2a = -10 so the focal length is a = -5 units... so the vertex is (4, 3 - (-5)) so vertex is at (4, 8) the general form of the equation I use is \[(x^2 - h) = 4a(y - k)\] \[(x - 4)^2 = 4\times(-5)(y - 8)\] just make y the subject for your solution . hope it helps